Please show all your work.

Your firm has total costs equal to 40Q1. Firm 2 has total costs of 50Q2. The demand you face is P=4000-5Q. You are the Stackelberg leader and firm 2 is the

Stackelberg follower. What is your output and profits?
Answer #1

The price curve for the product is P = 4000 - 5(Q1 + Q2)

Hence, profit curve for firm 1 = F1 = P*Q1 - C(Q1) = (4000 - 5 (Q1 + Q2))*Q1 - 40Q1 = 3660Q1 - 5Q12 - 5Q1Q2

Similarly, profit curve for firm 2 = F2 = P*Q2 - C(Q2) = (4000 - 5 (Q1 + Q2)) - 50Q2 = 3650Q1 - 5Q12 - 5Q1Q2

Hence, reaction curve for firm 1 and firm 2 are;

0 001 1F 3660 10Q1-5Q2

0 Q2 F2 = 3650-1002-5Q1 2aQ2

Now, firm 1 is being the leader. Hence, substituting firm 2's reaction function in firm 1's profit function, we have;

(3650 - 5Q1) 30Q12 22010Q1

Maximizinf F1, we have;

60Q122010 dQ1

Hence, Firm 1's output = 22010/60 = 366.83 = 367 say

Hence, Firm 1's profit is;

30Q12 22010Q1 24222005 3001 Fi